OpenStudy (anonymous):

the shortest side of a triangle is 7 inches shorter than the longest side. The middle side is twice as long as the short side. Find the length of all three sides if the perimeter is 31 inches.

OpenStudy (anonymous):

short side= x-7 mid= 2(x-7) long=x. (x-7)+2(x-7)+x=31. x-7+2x-14+x=31. 4x-21=31. 4x=52. x=13.

OpenStudy (radar):

Let a = length of long side and b = length of short side. c= lenght of remaining side. The following relationships are given. b=a-7 c=2b a+b+c = 31 inches. a+(a-7)+2(a-7)=31 a+a-7+2a-14=31 4a=31+7+14=52 a=13 b=6 c=12

OpenStudy (anonymous):

Long side= 13 in. short side= 6 in. middle side = 12. 13+6+12=31